The Toyota Prius is a compact hybrid hatchback. It is half the weight of the Escalade (1.3 tons ) has a 1.5 liter, 4 cylinder gasoline engine working together with an electric motor to produce power at a maximum rate of 76 horsepower. Its new EPA estimated combines highway and city mileage is 46 mpg. If you drive it for 15,000 miles over the course of a year, how much fuel energy is consumed?
1
Expert's answer
2020-04-26T18:52:43-0400

The Prius can make 46 miles consuming 1 gallon of fuel, or mathematically


"\frac{\text{Distance}}{\text{Fuel Volume}}=46."

Express Fuel Volume by multiplying both parts of the equation above by Fuel Volume and dividing both parts by 46:


"\text{Fuel Volume}=\frac{\text{Distance}}{46}=\frac{15000}{46}=326\text{ gallons}."

One gallon of gasoline is equivalent to approximately 130 800 000 Joules of energy, therefore, every 15000 km the Prius consumes


"326\cdot130800000=4.47\cdot10^{10}\text{ J}"


of fuel energy.